The manufacturer of a brand of juice drink claims that the mean content of its product is 210 ml with a standard deviation of 6.7 ml.

What is the probability that a randomly selected sample of the product contains more than or equal to 213 ml?

What is the probability that a sample contains less than 200 ml? 


1
Expert's answer
2021-12-20T10:27:59-0500
"z_1=\frac{213-210}{6.7}=0.45\\
P(z>0.45)=0.3264\\
z_2=\frac{200-210}{6.7}=-1.49\\P(z<-1.49)=0.0681"